a quantitative data set of 100 has a mean 25 and standard deviation of 4. At least how many observations lie between 17 and 33?
(round up to the nearest whole number)
Answer #1

Mean = = 25

SD = = 4

To find P(17 < X < 33):

k = (17 - 25)/4 = - 2

k = (33 - 25)/4 = 2

By Chebyshev's Theorem:

So,

Answer is:
At least number of observations lie between 17 and 33 is:given by:

100 X 0.75 = 75

So,

Answer is:

75